# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ Appendix D Deferred Proofs
D.8 Completion of the Proof in Section C.2
Let $B$ be the set of bid rates and interim allocation probabilities $(\beta_{i},p_{i})$ that satisfy Border’s Criterion. In Section C.2, we showed that the set of $(\beta_{i},p_{i})\in B$ satisfying (9) with strict inequality for $I\subsetneq[n]$ are induced by allocation probabilities. We now give a routine topological argument to show that we can extend this proof to all of $B$.
is compact being the inverse image of a closed set under a continuous mapping, and the set of $(\beta_{i},p_{i})$ that are induced by allocation probabilities is just the projection of this set onto the first two coordinates. Hence, it suffices to show that the set of $(\beta_{i},p_{i})\in B$ satisfying (9) with strict inequality for $I\subseteq[n]$ is dense in $B$. Given any such $(\beta_{i},p_{i})\in B$, letting $\epsilon>0$, define $(\beta_{i}^{\prime},p_{i}^{\prime})$ by
$\beta_{i}^{\prime}$ $=\beta_{i}+\epsilon$
Then, $(\beta_{i}^{\prime},p_{i}^{\prime})$ can be made arbitrarily close to $(\beta_{i},p_{i})$. They can be seen to satisfy Border’s Criterion where (9) is strict for $I\subsetneq[n]$ as follows. We have increased $\beta_{i}$ and decreased $p_{i}$ to get $(\beta_{i}^{\prime},p_{i}^{\prime})$ in such a way that both the left-hand side and right-hand side of (8) increase in the same amount to maintain equality. In (9) for $I\subsetneq[n]$, considering the changes caused by each coordinate $i\in I$ one at a time, the left-hand side increases by the same amount as in (8) but the right-hand side increases by a larger amount since the partial derivative of the right-hand side with respect to some $\beta_{i}\in I$ is (strictly) decreasing in $I$ (when all $\beta_{j}>0$ for $j\in I$). Therefore, with the $(\beta_{i}^{\prime},p_{i}^{\prime})$, (9) holds with strict inequality.

### v2

@@ E.4 Deferred Proofs from section˜C
E.9 Deferred Proofs from section˜6
Proof 26 (Proof of Lemma˜6.2).
The runtime claim is easily checked.
Let $\mathcal{S}_{i}^{(k)}=\{S\in\mathcal{S}^{(k)}:i\in S\}$. By the Chernoff bound, for any iteration $k$, agent $i$, and $0<\varepsilon<1$,
(E.9) $\Pr_{\mathcal{S}^{(k)}\sim\mu^{C}}\left(|\mathcal{S}_{i}^{(k)}|\leq(1-\varepsilon)\alpha_{i}C\right)\leq\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right).$
By the Hoeffding bound, for any fixed allocation probabilities $p_{i}^{S}$ and a fixed iteration $k$,
$\begin{split}\Pr_{\mathcal{S}^{(k)}\sim\mu^{C}}&\left(\left|\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in|\mathcal{S}_{i}^{(k)}|}p_{i}^{S}-\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}\mid i\in S]\right|\geq\delta\,\middle|\,|\mathcal{S}_{i}^{(k)}|\right)\leq 2\exp\left(-2|\mathcal{S}_{i}^{(k)}|\delta^{2}\right).\end{split}$
Therefore,
$\Pr_{\mathcal{S}^{(k)}\sim\mu^{C}}\left(\left|\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in|\mathcal{S}_{i}^{(k)}|}p_{i}^{S}-\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}\mid i\in S]\right|\geq\delta\right)\leq 2\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{2})+\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right).$
Let $E_{1}$ be the event where no LP at any iteration is infeasible. Using the allocation probabilities $p_{i}^{S}$ that satisfy (6.1) and (6.2) in the above equation, and the union bound over iterations $k\in[K]$ and agents $i\in[n]$, the above implies that
(E.10) $\Pr(E_{1}^{c})\leq\sum_{i=1}^{n}\left(2K\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{2})+K\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right)\right).$
By assumption, each $(p_{i}^{S^{*}})^{(k)}$ for $k\in[K]$ satisfies $\mathcal{F}(S^{*})$. Since $\mathcal{F}(S^{*})$ just consists of linear constraints, the average $\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S^{*}})^{(k)}$ satisfies $\mathcal{F}(S^{*})$. Algorithm˜2 sets $(p_{i}^{S^{*}})=\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S^{*}})^{(k)}$ on $E_{1}$. We now show that this output approximately satisfies (6.4).
On $E_{1}$, the multiplicative weights guarantee for solving linear programs, Theorem˜6.1, says for every $i$,
(E.11) $\left|\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}(p_{i}^{S})^{(k)}-\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)\right|\leq\delta+\left(\eta+\frac{\ln 2n}{\eta K}\right).$
For a fixed iteration $k$ and a randomly chosen $S$, $(p_{i}^{S})^{(k)}$ is a random variable whose law only depends on $y^{(k)}$ and $S$. Since $\mathcal{S}^{(k)}$ is independent of $y^{(k)}$, conditioning on $y^{(k)}$, the values of $(p_{i}^{S})$ for $S\in\mathcal{S}^{(k)}$ are just i.i.d. samples from the distribution of the $(p_{i}^{S})^{(k)}$ for $S\sim\mu$. This is a distribution over $[0,1]$ with mean $\operatorname*{\mathbb{E}}_{S\sim\mu}[(p_{i}^{S})^{(k)}\mid y^{(k)}]=\operatorname*{\mathbb{E}}_{S\sim\mu}[(p_{i}^{S})^{(k)}\mid(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]}]$.
Therefore, by Hoeffding’s inequality, for any $\delta^{\prime}>0$ and fixed $i$,
$\begin{split}\Pr_{\mathcal{S}^{(k)}\sim\mu^{C}}\left(\left|\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}(p_{i}^{S})^{(k)}-\operatorname*{\mathbb{E}}_{S\sim\mu}\left[(p_{i}^{S})^{(k)}\mid i\in S,(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]},|\mathcal{S}_{i}^{(k)}|\right]\right|\geq\delta^{\prime}\,\middle|\,|\mathcal{S}_{i}^{(k)}|,y^{(k)}\right)\\ \leq 2\exp(-2|\mathcal{S}_{i}^{(k)}|\delta^{\prime 2}).\end{split}$
By (E.9), this implies, for fixed $i$,
$\begin{split}\Pr_{\mathcal{S}^{(k)}\sim\mu^{C}}\left(\left|\frac{1}{|\mathcal{S}_{i}^{(k)}|}\sum_{S\in\mathcal{S}_{i}^{(k)}}(p_{i}^{S})^{(k)}-\operatorname*{\mathbb{E}}_{S\sim\mu}\left[(p_{i}^{S})^{(k)}\mid i\in S,(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]}\right]\right|\geq\delta^{\prime}\,\middle|\,(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]}\right)\\ \leq 2\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{\prime 2})+\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right).\end{split}$
Let $E_{2}$ be the event that the above does not happen for any iteration $k$ and any agent $i$, so by the union bound over iterations $k\in[K]$ and agents $i\in[n]$ and the above,
$\Pr(E_{2}^{c})\leq\sum_{i=1}^{n}\left(2K\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{\prime 2})+K\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right)\right)$
Let $E=E_{1}\cap E_{2}$. On $E$, by (E.11), the definition of $E_{2}$, and the triangle inequality,
$\begin{split}\left|\operatorname*{\mathbb{E}}_{\begin{subarray}{c}S\sim\mu\end{subarray}}\left[\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S})^{(k)}\,\middle|\,i\in S,(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]}\right]-\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)\right|\leq\delta^{\prime}+\delta+\left(\eta+\frac{\ln 2n}{\eta K}\right).\end{split}$
By the union bound,
$\Pr(E^{c})\leq\sum_{i=1}^{n}\left(2K\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{2})+2K\exp(-2(1-\varepsilon)\alpha_{i}C\delta^{\prime 2})+2K\exp\left(-\frac{\varepsilon^{2}C\alpha_{i}}{2}\right)\right).$
Substituting the choice of parameters in (6.9) along with $\varepsilon=\min\left\{\frac{1}{2},\sqrt{\frac{6\ln nK}{\underline{\alpha}C}}\right\}$ and $\delta^{\prime}=\delta=\sqrt{\frac{3\ln nK}{\underline{\alpha}C}}$ in the above two displays, we obtain the lemma statement.
Proof 27 (Proof of Footnote˜4).
Conditioned on the initially computed samples $\mathcal{S}^{(k)}$ and on the sets $S^{t}$ bidding at each time, at any time $t$, if $S^{t}=S$, agent $i$ wins the item with probability
(E.12) $p_{i}^{S}:=\operatorname*{\mathbb{E}}_{k\sim[K]}\left[(p_{i}^{S})^{(k)}\mid(\mathcal{S}^{(k^{\prime})})_{k^{\prime}\in[K]}\right]=\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S})^{(k)}.$
Thus, running Algorithm˜3 is equivalent to running BRB with preset allocation probabilities $p_{i}^{S}$. Since $\frac{1}{K}\sum_{k=1}^{K}(p_{i}^{S})^{(k)}$ is exactly the output of Algorithm˜2 when run with bidding agents $S$, we can use Lemma˜6.2 to see that
$\left|\operatorname*{\mathbb{E}}_{\begin{subarray}{c}S\sim\mu\end{subarray}}\left[p_{i}^{S}\,\middle|\,i\in S,(\mathcal{S}^{(k)})_{k\in[K]}\right]-\left(1-\prod_{j\in[n]}(1-\alpha_{j})\right)\right|\leq O\left(\sqrt{\frac{\log n}{K}}\right)\leq O\left(\sqrt{\frac{\log T}{T}}\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})\right)$
with probability at least $1-O(1/n^{2}T^{2})\geq 1-O(\operatorname{poly}(n,(1/\alpha_{j})_{j\in[n]})/T^{2})$. We obtain the high probability equilibrium utility guarantee by applying Theorem˜E.5.
The robustness guarantee follows from the fact that the each $(p_{i}^{S})^{(k)}$ satisfies $\mathcal{F}(S)$ and that $\mathcal{F}(S)$ is convex, so the expectation of the $(p_{i}^{S})^{(k)}$ satisfies $\mathcal{F}(S)$, which suffices to get the robustness guarantee from Lemma˜5.1.